OpenStudy (anonymous):

find the extreme on a closed interval f(x)=[-2,2] f'(x)=(-2x^2+2)/((x^2+1)^2)

OpenStudy (anonymous):

I mean f(x)=(2x)/(x^2+1)

myininaya (myininaya):

what is f?

myininaya (myininaya):

oh lol

myininaya (myininaya):

\[f'(x)=\frac{2(x^2+1)-2x(2x)}{(x^2+1)^2}=\frac{-2x^2+2}{(x^2+1)^2}\] good work on f' :)

myininaya (myininaya):

so x^2+1 is never 0

OpenStudy (anonymous):

thanks

myininaya (myininaya):

is the top ever 0?

myininaya (myininaya):

\[-2x^2+2=0 => 2=2x^2 => 1=x^2\]

myininaya (myininaya):

\[x=1,-1\]

OpenStudy (anonymous):

no

myininaya (myininaya):

why do you think that?

OpenStudy (anonymous):

well because you can sub nothing for x that would make it zero since it has +2

myininaya (myininaya):

so x=1 or x=-1 does not work?

OpenStudy (anonymous):

no you asked if the top could equal zero

myininaya (myininaya):

yes it can be zero when x=1,-1

OpenStudy (anonymous):

just solve f'(X)=0 and plug in the solutions

OpenStudy (anonymous):

so whats 0=(-2x^2+2)/((x^2+1)^2)

myininaya (myininaya):

\[-2x^2+2=0\] so when you solve this you get 2 real solutions

myininaya (myininaya):

\[x^2-1=0\] does look better to you?

myininaya (myininaya):

\[-2(x^2-1)=0 => x^2-1=0\]

OpenStudy (anonymous):

the deno is always positive, drop it. and we can get -2x^2+2=0

myininaya (myininaya):

\[x=\pm 1\]

OpenStudy (anonymous):

So -2(x^2-1)=0 so -2=0 which is none but wouldnt -1 squared be imaginary?

myininaya (myininaya):

\[(-1)^2=1 \]

myininaya (myininaya):

you are thinking of the square root of -1

OpenStudy (anonymous):

oh my bad, it's positive 1 not negative

myininaya (myininaya):

\[x^2-1=0 => x^2 =1 => x= \pm 1=x=1,-1\]

OpenStudy (anonymous):

why will it be negative 1 and +1?

myininaya (myininaya):

\[x^2-1=0\] \[(x-1)(x+1)=0\] \[x-1=0 \text{ or } x+1=0\] \[x=1 \text{ or } x=-1\]

OpenStudy (anonymous):

that's right, sorry for the misunderstanding.

myininaya (myininaya):

if you don;t like that way either lets go back to the first \[x^2-1=0 =>x^2=1 => \sqrt{x^2}= \sqrt{1} => |x|= 1 => x=\pm 1\]

myininaya (myininaya):

so anyways now plug in x= -2,-1,1,2 into f to get the value f(x)

OpenStudy (anonymous):

yeah I got that part, can you assist me on another problem pls?